OpenStudy (anonymous):

Solve the equation. 4y + 5 = – 31

OpenStudy (anonymous):

1-subtract 5 from each side 2-divide by 4 in each side

OpenStudy (anonymous):

y=-9

OpenStudy (anonymous):

okay thanks what about this one 3x – 7 = 5x + 19

OpenStudy (jiteshmeghwal9):

4y+5=-31 4y=-31-5 4y=-36 y=-36/4 =-9

OpenStudy (anonymous):

x=-13

OpenStudy (jiteshmeghwal9):

3x-7=5x+19 3x-5x=19+7 -2x=26 x=26/-2 x=-13.

OpenStudy (jiteshmeghwal9):

gt it?

OpenStudy (jiteshmeghwal9):

then give me a medal if u gt the question hahaha:)

OpenStudy (anonymous):

lol thank you so much!!!

OpenStudy (jiteshmeghwal9):

yw:)